TheSmartOne (thesmartone):

I kinda forgot how to do this, but I know it has to do with logarithms 17^(5x) = 8^(x+5) Solving for x.

TheSmartOne (thesmartone):

\(\Large 17^{5x} = 8^{x+5}\)

TheSmartOne (thesmartone):

If I take the log on both sides \(\Large 5x log(17) = (x+5)log8\) I don't really remember how to isolate x :P

TheSmartOne (thesmartone):

This is supposed to be easy, I remember solving questions like this, but I guess I'm getting rusty xD

jimthompson5910 (jim_thompson5910):

distribute the `log(8)` through and then get all the x terms to one side

TheSmartOne (thesmartone):

\(\Large 5x log(17) = xlog(8) + 5log(8)\) \(\Large 5x log(17) -xlog(8)= 5log(8)\) \(\Large x(5log(17) -log(8))= 5log(8)\) \(\Large x = \frac{5log(8)}{5log(17) -log(8)}\)

jimthompson5910 (jim_thompson5910):

you nailed it

TheSmartOne (thesmartone):

It was just that distributing that log(8) that had me stuck >.< Thanks, Jim! You're the best! :)

OpenStudy (mjdennis):

log(8) is just another constant, like "3" or "√2". Hope that helps in the future!

TheSmartOne (thesmartone):

mhmmm log(8) = log(2^3) = 3log(2)

OpenStudy (hedgepig):

I blame the obsolescence of slide rules for everyone's struggles with logarithmic properties.

OpenStudy (welshfella):

lol - that also confuses me sometimes !

OpenStudy (welshfella):

- distributing the log I mean.

TheSmartOne (thesmartone):

Yeah, I had a similar question for this assessment I had to take to show that I would be ready to start Calc in fall for Univ. But I was like this question has me stumped, and I knew how to solve them ages ago.

OpenStudy (hedgepig):

You really should review algebra in this case. Also your trig identities. Forget to do it, and you'll regret getting those annoying B's on your tests ;)

OpenStudy (welshfella):

Yes, however human error can't be eliminated.

OpenStudy (hedgepig):

That is true. It's nice though in calculus because generally you can take the inverse operation to check your answer. I easily got from a high 80 or low 90 to a high 90 just by taking time after a test to check my work. The more different the checking method is from the original, the more errors I tend to catch.

OpenStudy (welshfella):

Good.

OpenStudy (yumyum247):